two trains are 200 miles apart, and thier speeds differ by 10 mph. Find the speed of each train if they are traveling toward each other and will meet in 2 hours.
Answer by checkley77(12569)

D=RT
200=(R+R+10)2
200=(2R+10)2
200=4R+20
4R=200-20
4R=180
R=180/4
R=45 MPH. FOR THE SLOWER TRAIN.
45+10=55 MPH. FOR THE FASTER TRAIN.
PROO F:
200=(45+55)2
200=100*2
200=200